OpenStudy (anonymous):

Simplify. (3/8) ^3 * (40/9) ^2

OpenStudy (nurali):

\[(\frac{ 3 }{ 8 })^{3}*(\frac{ 40 }{ 9 })^{2}\] \[(\frac{ 27 }{ 512 })*(\frac{ 1600 }{ 81 })\]

OpenStudy (anonymous):

so do you flip the second fraction?

OpenStudy (anonymous):

No, only flip the second fraction if it's division. You flip the 2nd fraction in division to turn it into multiplication.

OpenStudy (anonymous):

so its 43200/41472

OpenStudy (nurali):

\[\frac{ 43200 }{ 41472 }\]

OpenStudy (nurali):

\[\frac{ 25 }{ 24 }\]

OpenStudy (anonymous):

how would you simply it

OpenStudy (anonymous):

He just divide 43200/41472 to get \[\frac{ 25 }{ 24 }\]

OpenStudy (anonymous):

It's in simplest form.